The length of a rectangle is 3 inches less than 4 times its width. The area of the rectangle is 7 sq. inches. Find the dimensions of the rectangle

Answer by TimothyLamb(4379)

---
y = 4x - 3
xy = 7
x(4x - 3) = 7
4xx - 3x - 7 = 0
--
the above quadratic equation is in standard form, with a=4, b=-3, and c=-7
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
4 -3 -7
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 1.75
x = -1
---
the negative root doesn't fit the problem statement, so use the positive root
---
x = 1.75 in
y = 4(1.75) - 3
y = 4 in
---
answer:
x = width = 1.75 in
y = length = 4 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php